# Discovery Prospects for a Minimal Dark Matter Model at Cosmic and Intensity Frontier Experiments

### v1

@@ 1 Introduction
Understanding the nature of dark matter (DM) is one of the most pressing challenges in particle physics. Cosmological and astrophysical observations provide compelling evidence for its existence, yet its fundamental nature remains unknown. The absence of a discovery of weakly interacting massive particle (WIMP) dark matter has strongly motivated the exploration of alternative dark matter candidates. In particular, scenarios featuring dark matter with a mass below the electroweak scale and residing in its own hidden sector have gained significant attention Abercrombie:2015wmb ; Alexander:2016aln .
@@
Dark photons are a possible minimal extension to the Standard Model (SM) that can provide a portal to a larger hidden sector Ackerman:2008kmp ; Feng:2008mu ; Chu:2011be . In this paper, we consider a minimal secluded dark matter model, where dark matter is a Dirac fermion, $\chi$, charged only under a dark $U(1)$ gauge symmetry Pospelov:2007mp . The dominant interaction between the dark matter and the SM is mediated by a kinetically mixed dark photon, $Z_{D}$, with mass $m_{Z_{D}}<m_{\chi}$. In this regime, the dark photon decays into SM particles, making it an important target for visibly decaying long-lived particle searches at beam-dump experiments, LHCb, Belle II, and dedicated LHC auxiliary detectors Antel:2023hkf .
@@
In this scenario, the dark matter relic abundance is determined by two key processes: direct dark matter production from the SM, $f\bar{f}\to\chi\bar{\chi}$, and dark matter annihilation into dark photons, $\chi\bar{\chi}\to Z_{D}Z_{D}$, which in turn decay back to the SM Bernal:2015ova ; Krnjaic:2017tio ; Berlin:2017ftj ; Evans:2017kti ; Hambye:2019dwd ; Du:2020avz ; Tapadar:2021kgw ; Hryczuk:2021qtz ; Slatyer:2015jla ; DelNobile:2015uua ; Fernandez:2021iti . Above a certain value of the kinetic mixing parameter, $\epsilon$, the hidden sector reaches thermal equilibrium with the SM in the early universe, and the dark matter abundance is controlled only by the dark gauge coupling, $\alpha_{D}$. We refer to this value as the thermalization floor. Below the thermalization floor, dark matter is not in equilibrium with the SM, requiring a full solution of coupled Boltzmann equations for the hidden sector temperature as well as the number of dark matter particles Chu:2011be . The observed dark matter density can be realized along a continuous curve in the $\alpha_{D}$-$\epsilon$ plane, which interpolates between a freeze-in mechanism at small $\alpha_{D}$ and a secluded freeze-out scenario at larger values of $\alpha_{D}$.
@@
As we will demonstrate, the parameter space where the hidden sector was in thermal equilibrium with the SM at the time of dark matter freeze-out—the so-called WIMP-next-door regime Evans:2017kti —is now entirely excluded by a combination of direct and indirect detection constraints. However, large regions of parameter space of the model remain viable, particularly those where dark matter attains its relic abundance via the freeze-in mechanism or through freeze-out processes occurring while the hidden sector is undergoing non-adiabatic evolution. This presents exciting opportunities: a dark photon discovery at future beam-dump experiments such as DUNE Berryman:2019dme ; DUNE:2020ypp , SHiP SHiP:2021nfo ; Alekhin:2015byh , and DarkQuest Berlin:2018pwi ; Apyan:2022tsd would be fully consistent with this minimal dark matter scenario, even in the continued absence of signals in direct detection experiments. Additionally, while much of the viable parameter space predicts dark matter-nuclear scattering cross-sections that lie below the so-called neutrino fog, and is therefore challenging for direct detection experiments to probe, we find there still exist regions of parameter space where this model can be uniquely tested through future direct detection efforts.
@@ 2.2 Thermal History
[…]
Here the matrix element |¯$\left|\overline{\mathcal{M}}\right|^{2}$ is averaged over initial state spin and summed over final state spins. In App. A, we present detailed expressions for $A(s)$ for both processes shown in Fig. 1.
@@
[…]
@@
[…]
@@
$\mathcal{C}^{n}_{\chi\bar{\chi}\to Z_{D}Z_{D}}(\tilde{T})=$ $\,n_{\mathrm{eq}}^{2}(\tilde{T})\langle\sigma v\rangle_{Z_{D}Z_{D}\to\chi\bar{\chi}}$ (13)
@@
In solving these equations, we write $\rho_{HS}=\rho_{Z_{D}}+2\rho_{\chi}$ where the factor of 2 accounts for the contributions of $\chi$ and $\bar{\chi}$. We take the energy density and pressure of the $Z_{D}$ to be given by the equilibrium distribution for a relativistic boson, $\rho_{Z_{D}}=3P_{Z_{D}}=g_{Z_{D}}(\frac{\pi^{2}}{30})\tilde{T}^{4}$. For the dark matter, in order to incorporate the contribution from its rest energy when non-relativistic, we adopt the Maxwell-Boltzmann results
@@
We use $x\equiv m_{\chi}/T$ as our independent variable and rewrite Eq. 14 to solve for $\tilde{T}$ and the DM yield $Y_{\chi}\equiv n_{\chi}/s(T)$, where $s(T)=g_{*S,SM}T^{3}$ is the SM entropy density. Using the conservation of SM entropy, we translate a time derivative into a derivative with respect to $x$ as
@@
$\frac{d\tilde{T}}{dx}$ $=\Bigg{(}s^{2}\sum_{f}\langle\sigma vE\rangle_{f\bar{f}\rightarrow\chi\bar{\chi}}Y_{f}^{2}(T)-s^{2}\sum_{f}\langle\sigma vE\rangle_{\chi\bar{\chi}\rightarrow f\bar{f}}Y_{\chi}^{2}-4H\rho_{Z_{D}}-2sAB\frac{dY_{\chi}}{dx}-6HsY_{\chi}\tilde{T}\Bigg{)}\times$ (17)
$\left(\frac{d\rho_{Z_{D}}}{d\tilde{T}}-2\frac{m_{\chi}}{\tilde{T}^{2}}\frac{dB}{d\tilde{T}}sY_{\chi}\right)^{-1}$
@@ 2.3 Numerical solutions
There are four distinct regions along each mass contour. At small $\alpha_{D}$, the curves start out in the “freeze-in” regime such that DM abundance is determined only by direct DM production from the SM, which is proportional to $\epsilon^{2}\alpha_{D}$. As $\alpha_{D}$ becomes larger, the process $\chi\bar{\chi}\to Z_{D}Z_{D}$ starts to appreciably deplete DM, which means that more DM (and thus more energy) must be provided by out-of-equilibrium production from the SM to obtain the right DM abundance. In this regime, the curves in Fig. 2 require $\epsilon$ to increase with increasing $\alpha_{D}$. The dynamics in this regime result from a nontrivial interplay of out-of-equilibrium processes with equilibrium processes within the hidden sector, giving rise to steep “reannihilation” Chu:2011be ; Berger:2018xyd and more shallowly sloped “leak-in” Evans:2019vxr phases. We refer to these two phases collectively as “out-of-equilibrium freeze-out”. Finally, at sufficiently large values of $\epsilon$, the HS equilibrates with the SM before DM freeze-out. In this “WIMP-next-door” regime, the DM relic abundance becomes independent of $\epsilon$. The minimum value of $\epsilon$ that allows for this equilibration, as a function of $m_{\chi}$, defines a thermalization floor [^2] These results update and correct the vector portal results of Evans:2017kti , which did not account for the in-medium suppression of the kinetic mixing in evaluating the thermalization floor..
@@
It will be convenient in what follows to split up the predictions of this model into three distinct regimes: (i) freeze-in, (ii) out-of-equilibrium freeze-out, and (iii) above the thermalization floor. Quantitatively, we identify the thermalization floor $\alpha_{D}^{TF}$ for each DM mass by finding the value of $\alpha_{D}$ at which the relic abundance curve becomes $\epsilon$-independent. We separate the freeze-in regime from the out-of-equilibrium freeze-out regime using the minimum value of $\epsilon$ on the relic curve in Fig. 2, so that the freeze-in branch refers to the portion of the relic curve with $\alpha_{D}<\alpha_{D}(\epsilon_{\mathrm{min}})$. Thus the out-of-equilibrium freeze-out regime is characterized by $\alpha_{D}(\epsilon_{\mathrm{min}})<\alpha_{D}<\alpha_{D}^{TF}$. [^3] Dark photons with $\epsilon<\epsilon_{\mathrm{min}}$ can still be consistent with this minimal model of DM if the hidden sector is populated by some non-SM source in the early universe Arvanitaki:2021qlj ; Fernandez:2021iti . Here we focus on the more predictive scenario where interactions with the SM dominate the production.
@@
Figure 3: The thermalization floor (red) and minimum mixing $\epsilon_{\text{min}}$ separating freeze-in and freeze-out regimes (blue) mapped into the $m_{Z_{D}}$-$\epsilon$ plane, for fixed $m_{Z_{D}}/m_{\chi}=10$ (darker) and $m_{Z_{D}}/m_{\chi}=100$ (lighter). Additional curves showing excluded and projected reach from accelerator-based experiments are taken from Ref. Batell:2022dpx ; SHiP:2021nfo .
@@
In Fig. 3, we show how the thermalization floor and the minimum-mixing line $\epsilon_{\mathrm{min}}$ can be translated into the dark photon parameter space. The thermalization floor and minimum-mixing line are defined as a function of DM mass. To map them into the dark photon parameter space, we consider fixed values of the mass ratio $m_{Z_{D}}/m_{\chi}$. For illustrative purposes, we show results for $m_{\chi}=10m_{Z_{D}}$ and $m_{\chi}=100m_{Z_{D}}$. As we can see from the figure, both the QCD phase transition and DM production from $Z$ decays give the thermalization floor and the minimum-mixing a notable $m_{\chi}$ dependance, in contrast to the simpler $m_{\chi}$ dependence exhibited away from mass thresholds Evans:2019vxr ; Fernandez:2021iti . In Sec. 3 we will discuss in detail the interplay between searches for the dark photon and constraints on DM from direct and indirect detection.
@@ 3.1 Indirect Detection
Since dark photons decay promptly into pairs of SM particles, the products of DM annihilations are energetic SM particles, which can leave imprints in astrophysical and cosmological datasets. We focus here on the constraints on DM annihilation provided by the anisotropy spectrum of the cosmic microwave background (CMB). Fermi-LAT observations of dwarf galaxies Fermi-LAT:2015att ; Fermi-LAT:2016uux as well as AMS-02 observations of positron spectra AMS:2013fma can provide similar constraints on the DM annihilation cross-section and can be nominally more constraining for DM masses in the range of primary interest to us ($5\,\mathrm{GeV}\lesssim m_{\chi}\lesssim 100$ GeV). However, searches for DM annihilation products from galaxies today are subject to sizable systematic uncertainties relating to the modeling of DM haloes and astrophysical sources of cosmic rays. CMB constraints, on the other hand, primarily depend on the properties of the universe at times when it was well-described by perturbation theory.
@@
Figure 4: Indirect detection constraints from CMB on the ($m_{Z_{D}},\epsilon$) plane for several values of $m_{\chi}$ plotted with current and projected beam dump limits and other accelerator based experiments as LHCb and Belle II. The CMB constraints are not monotonic in behavior as a function of the DM mass because of the Sommerfeld enhancements at larger masses. Additional curves showing excluded and projected reach from accelerator-based experiments are taken from Ref. Batell:2022dpx ; SHiP:2021nfo .
Imposing that the observed relic abundance of DM is obtained fixes $\alpha_{D}$ as a function of $m_{\chi}$ and $m_{Z_{D}}$. Since we found this result to be approximately independent of the value $m_{Z_{D}}$ for dark photon masses sufficiently below $m_{\chi}$ (and $m_{Z}$), we take $\alpha_{D}$ to be a function only of the kinetic mixing parameter $\epsilon$ (for fixed $m_{\chi}$) and retain $m_{Z_{D}}$ dependence only as defined in equations Eqs. 2220. Only the out-of-equilibrium freeze-out and thermalized regions predict DM annihilation cross sections large enough to be constrained by indirect detection. We thus consider the portion of the relic curves with $\alpha_{D}(\epsilon)>\alpha_{D}(\epsilon_{\mathrm{min}})$, which determines a unique value of $\alpha_{D}$ at a given $\epsilon$. The resulting bounds are shown in Fig. 4 for several values of $m_{\chi}$. We see that, as expected, CMB constraints are strongest for the lightest DM masses, extending down to the smallest values of $\epsilon$ and thus to the smallest overall annihilation cross sections. For $m_{\chi}<15$ GeV, the right boundary of the exclusion is set by the requirement that $m_{Z_{D}}<m_{\chi}/10$, as this is the regime where our calculations are self-consistent. As $m_{\chi}$ increases, the value of the excluded annihilation cross-section decreases, and thus the CMB constraints become successively weaker for $15\,\mathrm{GeV}\lesssim m_{\chi}\lesssim 50$ GeV. Finally for $m_{\chi}=100$ GeV the excluded parameter space increases again, as the Sommerfeld enhancement begins to kick in. Note that sizable portions of the projected sensitivity regions for DarkQuest, SHiP, and DUNE remain unconstrained [^4] A similar conclusion holds for the FASER2 experiment Feng:2022inv , which is not shown in the figure..

### v2

@@ 1 Introduction
Understanding the nature of dark matter (DM) is one of the most pressing challenges in particle physics. Cosmological and astrophysical observations provide compelling evidence for its existence, yet its fundamental nature remains unknown. The absence of a discovery of weakly interacting massive particle (WIMP) dark matter has strongly motivated the exploration of alternative dark matter candidates. In particular, scenarios featuring dark matter with a mass below the electroweak scale and residing in its own hidden sector have gained significant attention Abercrombie:2015wmb ; Alexander:2016aln ; Battaglieri:2017aum .
@@
Dark photons are a possible minimal extension to the Standard Model (SM) that can provide a portal to a larger hidden sector Ackerman:2008kmp ; Feng:2008mu ; Chu:2011be ; Fabbrichesi:2020wbt . In this paper, we consider a minimal secluded dark matter model, where dark matter is a Dirac fermion, $\chi$, charged only under a dark $U(1)$ gauge symmetry Pospelov:2007mp . The dominant interaction between the dark matter and the SM is mediated by a kinetically mixed dark photon, $Z_{D}$, with mass $m_{Z_{D}}<m_{\chi}$. In this regime, the dark photon decays into SM particles, making it an important target for visibly decaying long-lived particle searches at beam-dump experiments, LHCb, Belle II, and dedicated LHC auxiliary detectors Antel:2023hkf .
@@
In this scenario, the dark matter relic abundance is determined by two key processes: direct dark matter production from the SM, $f\bar{f}\to\chi\bar{\chi}$, and dark matter annihilation into dark photons, $\chi\bar{\chi}\to Z_{D}Z_{D}$, which in turn decay back to the SM Bernal:2015ova ; Slatyer:2015jla ; DAgnolo:2015ujb ; DelNobile:2015uua ; Evans:2017kti ; Hambye:2019dwd ; Fernandez:2021iti ; Bhattiprolu:2023akk ; Boddy:2024vgt . Above a certain value of the kinetic mixing parameter, $\epsilon$, the relic abundance curves become independent of $\epsilon$, and the dark matter abundance is controlled only by the dark gauge coupling, $\alpha_{D}$. For simplicity we refer to this value of $\epsilon$ as the thermalization floor. Below the thermalization floor, the dark sector temperature evolves non-adiabatically, requiring a full solution of coupled Boltzmann equations for the hidden sector temperature as well as the number of dark matter particles Chu:2011be . The observed dark matter density can be realized along a continuous curve in the $\alpha_{D}$-$\epsilon$ plane, which interpolates between a freeze-in mechanism at small $\alpha_{D}$ and a secluded freeze-out scenario at larger values of $\alpha_{D}$.
@@
The aim of this work is to provide an up-to-date assessment of the detection prospects of this minimal model. As we will demonstrate for the first time, the parameter space of this minimal model above the thermalization floor—the so-called WIMP-next-door regime Evans:2017kti —is now entirely excluded by a combination of direct and indirect detection constraints. However, large regions of parameter space of the model remain viable, particularly those where dark matter attains its relic abundance via the freeze-in mechanism or through freeze-out processes occurring while the hidden sector is undergoing non-adiabatic evolution. In this paper, we also highlight exciting opportunities for future experiments: as we demonstrate in this work, a dark photon discovery at future beam-dump experiments such as DUNE Berryman:2019dme ; DUNE:2020ypp , SHiP SHiP:2021nfo ; Alekhin:2015byh , and DarkQuest Berlin:2018pwi ; Apyan:2022tsd would be fully consistent with this minimal dark matter scenario, even in the continued absence of signals in direct detection experiments. Additionally, while much of the viable parameter space predicts dark matter-nuclear scattering cross-sections that lie below the so-called neutrino fog, and is therefore challenging for direct detection experiments to probe, we find there still exist regions of parameter space where this model can be uniquely tested through future direct detection efforts.
@@ 2.2 Thermal History
$\mathcal{A}(s)\equiv\int d\Pi_{f}(2\pi)^{4}\delta^{4}(\sum p)\left|\overline{\mathcal{M}}\right|^{2}.$ (10)
Here the matrix element $\left|\overline{\mathcal{M}}\right|^{2}$ is averaged over initial state spin and summed over final state spins. In App. A, we present detailed expressions for $A(s)$ for both processes shown in Fig. 1.
@@
$=$ $\int d\Pi_{i}|\overline{\mathcal{M}}_{f\bar{f}\rightarrow\chi\bar{\chi}}|^{2}(2\pi)^{4}\delta^{4}(\sum_{i}p_{i})f_{f}^{\text{eq}}(T)f_{\bar{f}}^{\text{eq}}(T)$
$=$ $\frac{g_{f}g_{\bar{f}}}{8\pi^{4}}\frac{T}{8}\int_{\max(4m_{\chi},4m_{f}^{2})}^{\infty}ds\sqrt{s-4m_{f}^{2}}\mathcal{A}_{f\bar{f}\rightarrow\chi\bar{\chi}}(s)K_{1}(\sqrt{s}/T),$
@@
$=$ $\int d\Pi_{i}|\overline{\mathcal{M}}_{f\bar{f}\rightarrow\chi\bar{\chi}}|^{2}(2\pi)^{4}\delta^{4}(\sum_{i}p_{i})f_{f}^{\text{eq}}(T)f_{\bar{f}}^{\text{eq}}(T)(E_{f}+E_{\bar{f}})$
$=$ $\frac{g_{f}g_{\bar{f}}}{8\pi^{4}}\frac{T}{8}\int_{\max(4m_{\chi},4m_{f}^{2})}^{\infty}ds\sqrt{s}\sqrt{s-4m_{f}^{2}}\mathcal{A}_{f\bar{f}\rightarrow\chi\bar{\chi}}(s)K_{2}(\sqrt{s}/T).$
@@
[…]
@@
In solving these equations, we write $\rho_{HS}=\rho_{Z_{D}}+2\rho_{\chi}$ where the factor of 2 accounts for the contributions of $\chi$ and $\bar{\chi}$. Recalling that our DM is Dirac, we will ultimately require $\rho_{\chi}$ to make up half the observed DM relic abundance. We take the energy density and pressure of the $Z_{D}$ to be given by the equilibrium distribution for a relativistic boson, $\rho_{Z_{D}}=3P_{Z_{D}}=g_{Z_{D}}(\frac{\pi^{2}}{30})\tilde{T}^{4}$. For the dark matter, in order to incorporate the contribution from its rest energy when non-relativistic, we adopt the Maxwell-Boltzmann results
@@
We use $x\equiv m_{\chi}/T$ as our independent variable and rewrite Eq. 14 to solve for $\tilde{T}$ and the DM yield $Y_{\chi}\equiv n_{\chi}/s(T)$, where $s(T)=\frac{2\pi^{2}}{45}g_{*S,SM}T^{3}$ is the SM entropy density. Using the conservation of SM entropy, we translate a time derivative into a derivative with respect to $x$ as
@@
[…]
$\left(A\frac{d\rho_{Z_{D}}}{d\tilde{T}}+2sY_{\chi}A\frac{dB}{d\tilde{T}}\right)^{-1}$
@@ 2.3 Numerical solutions
There are four distinct regions along each mass contour. At small $\alpha_{D}$, the curves start out in the “freeze-in” regime such that DM abundance is determined only by direct DM production from the SM, which is proportional to $\epsilon^{2}\alpha_{D}$. As $\alpha_{D}$ becomes larger, the process $\chi\bar{\chi}\to Z_{D}Z_{D}$ starts to appreciably deplete DM, which means that more DM (and thus more energy) must be provided by out-of-equilibrium production from the SM to obtain the right DM abundance. In this regime, the curves in Fig. 2 require $\epsilon$ to increase with increasing $\alpha_{D}$. The dynamics in this regime result from a nontrivial interplay of out-of-equilibrium processes with equilibrium processes within the hidden sector, giving rise to steep “reannihilation” Chu:2011be ; Berger:2018xyd and more shallowly sloped “leak-in” Evans:2019vxr phases. [^2] The characteristic out-of-equilibrium dynamics realized here rely on the existence of a dark radiation bath, which is the key motivation for our restriction $m_{Z_{D}}<m_{\chi}/10$. Adding an additional dark radiation species to furnish a “dark sink” Bhattiprolu:2023akk allows these out-of-equilibrium dynamics to be realized for less-restrictive choices of the dark photon mass. We refer to these two phases collectively as “out-of-equilibrium freeze-out”. Finally, at sufficiently large values of $\epsilon$, the HS equilibrates with the SM before DM freeze-out. In this “WIMP-next-door” regime, the DM relic abundance becomes independent of $\epsilon$. The minimum value of $\epsilon$ that allows for this equilibration, as a function of $m_{\chi}$, defines a thermalization floor [^3] These results update and correct the vector portal results of Evans:2017kti , which did not account for the in-medium suppression of the kinetic mixing in evaluating the thermalization floor..
@@
It will be convenient in what follows to split up the predictions of this model into three distinct regimes: (i) freeze-in, (ii) out-of-equilibrium freeze-out, and (iii) above the thermalization floor. Quantitatively, we identify the thermalization floor $\alpha_{D}^{TF}$ for each DM mass by finding the value of $\alpha_{D}$ at which the relic abundance curve becomes $\epsilon$-independent. It is worth bearing in mind a subtlety here: for values of $\epsilon$ immediately above this “thermalization floor,” the dark sector is not actually equilibrated with the SM throughout the freeze-out process. Rather, in this regime the energy injection from the SM into the dark sector after kinetic decoupling of the two sectors is negligible in comparison with the red-shifting of the HS energy bath, so that the HS temperature still evolves adiabatically.
We separate the freeze-in regime from the out-of-equilibrium freeze-out regime using the minimum value of $\epsilon$ on the relic curve in Fig. 2, so that the freeze-in branch refers to the portion of the relic curve with $\alpha_{D}<\alpha_{D}(\epsilon_{\mathrm{min}})$. Thus the out-of-equilibrium freeze-out regime is characterized by $\alpha_{D}(\epsilon_{\mathrm{min}})<\alpha_{D}<\alpha_{D}^{TF}$. [^4] Dark photons with $\epsilon<\epsilon_{\mathrm{min}}$ can still be consistent with this minimal model of DM if the hidden sector is populated by some non-SM source in the early universe Arvanitaki:2021qlj ; Fernandez:2021iti . Here we focus on the more predictive scenario where interactions with the SM dominate the production.
@@
Figure 3: The thermalization floor (red) and minimum mixing $\epsilon_{\text{min}}$ separating freeze-in and freeze-out regimes (blue) mapped into the $m_{Z_{D}}$-$\epsilon$ plane, for fixed $m_{\chi}/m_{Z_{D}}=10$ (darker) and $m_{\chi}/m_{Z_{D}}=100$ (lighter). Additional curves showing excluded regions (shaded gray) and projected reach (colored dotted lines) from accelerator-based experiments are taken from Refs. Batell:2022dpx ; SHiP:2021nfo , while exclusions from Supernova 1987a are taken from from Ref. Chang:2016ntp .
@@
In Fig. 3, we show how the thermalization floor and the minimum-mixing line $\epsilon_{\mathrm{min}}$ can be translated into the dark photon parameter space. The thermalization floor and minimum-mixing line are defined as a function of DM mass. To map them into the dark photon parameter space, we consider fixed values of the mass ratio $m_{Z_{D}}/m_{\chi}$. For illustrative purposes, we show results for $m_{\chi}=10m_{Z_{D}}$ and $m_{\chi}=100m_{Z_{D}}$. As we can see from Fig. 2 and Fig. 3, both the QCD phase transition and DM production from $Z$ decays give the thermalization floor and the minimum-mixing a notable $m_{\chi}$ dependance, in contrast to the simpler $m_{\chi}$ dependence exhibited away from mass thresholds Evans:2019vxr ; Fernandez:2021iti . In Sec. 3 we will discuss in detail the interplay between searches for the dark photon and constraints on DM from direct and indirect detection.
@@ 3.1 Indirect Detection
Since dark photons decay promptly into pairs of SM particles, the products of DM annihilations are energetic SM particles, which can leave imprints in astrophysical and cosmological datasets. We focus here on the constraints on DM annihilation provided by the anisotropy spectrum of the cosmic microwave background (CMB). Fermi-LAT observations of dwarf galaxies Fermi-LAT:2015att ; Fermi-LAT:2016uux as well as AMS-02 observations of positron spectra AMS:2013fma can provide similar constraints on the DM annihilation cross-section and can nominally contribute additional sensitivity for DM masses above $10$ GeV. However, searches for DM annihilation products from galaxies today are subject to sizable systematic uncertainties relating to the modeling of DM haloes and astrophysical sources of cosmic rays. CMB constraints, on the other hand, primarily depend on the properties of the universe at times when it was well-described by perturbation theory. In addition, as shown for this model in Evans:2017kti , the overall sensitivity of cosmic ray searches is broadly comparable to that from the CMB. For this reason, we do not add cosmic ray observations to our list of constraints.
@@
Figure 4: Indirect detection constraints from CMB on the ($m_{Z_{D}},\epsilon$) plane for several values of $m_{\chi}$ plotted with current and projected beam dump limits and other accelerator based experiments as LHCb and Belle II. The CMB constraints are not monotonic in behavior as a function of the DM mass because of the Sommerfeld enhancements at larger masses. Additional curves showing excluded regions (shaded gray) and projected reach (colored dotted lines) from accelerator-based experiments are taken from Refs. Batell:2022dpx ; SHiP:2021nfo , while exclusions from Supernova 1987a are taken from from Ref. Chang:2016ntp .
Imposing that the observed relic abundance of DM is obtained fixes $\alpha_{D}$ as a function of $m_{\chi}$ and $m_{Z_{D}}$. Since we found this result to be approximately independent of the value $m_{Z_{D}}$ for dark photon masses sufficiently below $m_{\chi}$ (and $m_{Z}$), we take $\alpha_{D}$ to be a function only of the kinetic mixing parameter $\epsilon$ (for fixed $m_{\chi}$) and retain $m_{Z_{D}}$ dependence only as defined in equations Eqs. 2220. Only the out-of-equilibrium freeze-out and thermalized regions predict DM annihilation cross sections large enough to be constrained by indirect detection. We thus consider the portion of the relic curves with $\alpha_{D}(\epsilon)>\alpha_{D}(\epsilon_{\mathrm{min}})$, which determines a unique value of $\alpha_{D}$ at a given $\epsilon$. The resulting bounds are shown in Fig. 4 for several values of $m_{\chi}$. We see that, as expected, CMB constraints are strongest for the lightest DM masses, extending down to the smallest values of $\epsilon$ and thus to the smallest overall annihilation cross sections. For $m_{\chi}<15$ GeV, the right boundary of the exclusion is set by the requirement that $m_{Z_{D}}<m_{\chi}/10$, as this is the regime where our calculations are self-consistent. As $m_{\chi}$ increases, the value of the excluded annihilation cross-section decreases, and thus the CMB constraints become successively weaker. Note that sizable portions of the projected sensitivity regions for DarkQuest, SHiP, and DUNE remain unconstrained [^5] A similar conclusion holds for the FASER2 experiment Feng:2022inv , which is not shown in the figure..